OpenStudy (anonymous):

simplify (3a^0b^2)(2a^-3b^2)^2

OpenStudy (anonymous):

you can start with the fact that \(a^0=1\) and eliminate it by writing \[(3b^2)(2a^{-3}b^2)^2\]

OpenStudy (anonymous):

then perhaps use \((b^n)^m=b^{mn}\) to change the second parentheses to \[3b^2(2^2a^{-6}b^4)\]

OpenStudy (anonymous):

then multiply out, adding the exponents as you go, get \[12a^{-6}b^6\]

OpenStudy (anonymous):

if you need to answer with positive exponents only, you can rewrite this one as \[\frac{12b^6}{a^6}\]

OpenStudy (anonymous):

so its it 12b^/a^6

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

yes yw

OpenStudy (anonymous):

how can u write this using rational exponents

OpenStudy (anonymous):



OpenStudy (anonymous):